ID |
Date |
Author |
Type |
Category |
Subject |
84
|
Tue Sep 2 22:48:47 2008 |
Dmass | Laser | PSL | Connectors needed |
I am trying to chase down the source of the long time scale sharp fluctiations in intensity. I need to start reading out the diagnostic PD from inside the MOPA head unit to see if the power fluctuations are coming before or after the amplification (e.g. NPRO or MOPA problems).
I need the following connectors from Wilson house: |
Attachment 1: connectors.png
|
|
83
|
Tue Sep 2 15:06:42 2008 |
Aidan | Laser | Fiber | Attempted calibration of fiber noise ... issues! |
I tried to calibrate the fiber phase noise measurements in DTT by determining the fringe visibility (peak-to-peak size) in counts (to determine radians per count). I shifted the carrier on the VCO by 100Hz to 200.0001MHz and left the other signal generator (LO for both mixers) at exactly 200MHz. The demodulated signals from the PDs then looked like nice 100Hz sine waves. It was only then that I noticed that the amplitude of these sine waves was varying by 50% over a timescale of 20 seconds to a number of minutes.
NPRO level: 1.500A
I plotted 20 minutes worth of Max-Min data from the 2 PDs in DataViewer (see attached pdf) - there was 100Hz frequency difference between the signal generators for all but the first four minutes or so of this plot. In principle this plot shows the fringe visibility in counts and can be used for calibration purposes. As you can see, though, the pk-pk value is fluctuating with time. Presumably, as nothing has been altered on the table over the last week, the fringe visibility has been fluctuating for all previous noise measurements. Therefore only a rough calibration is possible. So here it is ...
C2: OMS-SUS_TOP1_INI_65536: pi radians: [-2750, -700] counts -> ~1500 urad/count
C2: OMS-SUS_TOP2_INI_65536: pi radians: [-3600, 4000] counts -> ~410 urad/count
Eric G happened to come in shortly after I had discovered this. He suggested that the mode-matching between the two beams may not be good enough. I will look into this. |
Attachment 1: peak_peak_calibration.pdf
|
|
82
|
Tue Sep 2 11:01:00 2008 |
Aidan | Electronics | Fiber | Fiber stabilization loop diagram |
Attached is a diagram of the current version of the stabilization loop. The beam (red) into the AOM is the output from the fiber and the other red line represents the reference beam. |
81
|
Wed Aug 20 23:19:39 2008 |
Dmass | Laser | Doubling | Preliminary Doubling MZ Setup |
A prelim expermintal setup for measuring the phase/freq noise of 532 nm light generated by PPKTP crystals relative to the 1064 nm seed light from the YAG laser.
The mirrors in the Mach-Zehnder will all be dichroically reflective. The pink mirrors on the output will be reflective in one wavelength, and transmissive in the other.
I may need to add some sort of absorption/reflection before the PPKTP crystals for the green light if it is emitted in both directions by the SHG process.
I should be able to look at the difference between the MZ output in 1064 and 532 to isolate the more interesting sources of noise (phase/frequency) from the acoustic noise. |
Attachment 1: MachZenderPPKTP.PNG
|
|
80
|
Mon Aug 18 23:46:42 2008 |
Dmass | Computing | DAQ | More on the DAQ |
It appeared today that the DAQ channels wouldnt record. Only one would write no matter how I edited the .ini file.
I rebuilt everything, and the channels had the _32xxx suffix instead of the _DAQ suffix, and proceeded to write to the DAQ, allowing Masha to continue to slave away. |
79
|
Fri Aug 15 15:59:07 2008 |
Dmass | Computing | DAQ | DAC DAQ and ADC working |
The problem was the timing. I changed what the simulink .mdl file wanted for a clock signal to what it was actally getting (64 kHz), and everything was magically ok.
It appears that the ADC can function at 32kHz with a 64kHz clock signal, but the DAC apparently can not.
Several PC board were removed and/or inspected before I realized that I should change this. |
78
|
Fri Aug 15 02:17:46 2008 |
Dmass | Electronics | General | no DAC output voltage when running OMS didgital system |
Quote: | I tried to use the OMS digital system for a measurement, but I couldn't get the DAC output to work.
Symptoms:
- All EPICS buttons work, and digital read-backs make sense
- ADC works just fine
- DAC output voltage is zero, no matter what the filter module output says
- Watchdog is enabled (green) |
I talked to Alex, he suggested it was a hardware problem. I then talked to Rich, who in turn suggested it was a software problem.
As far as we know, nothing has changed about the hardware configuration since it was working earlier this year.
That said, I tried several things to troubleshoot.
I borrowed the SCSI breakout board from Wilson house, and plugged it into the anti aliasing box.
I was able to drive the AA box and see the signal make it through.
I turned everything in the crate off.
Then I moved the LASTI timing box below the blue PCIX box so I could open the latter up.
I grounded myself with a static strap, and the breakout board with a wire to the chassis of the PCIX box.
I then put the breakout board directly at the output of the DAC inside the blue PCIX box. I was
careful to elevate it via rubber feet and not let the bottom touch anything.
I then turned everything back on and restarted the fe machine with:
*ssh oms
*startoms
and from ws1:
*telnet fb0 8087
*shutdown
:noteworthy errors from the above follow at the end of this entry:
This got everything up and as working as it was before I started tinkering. I sent a booming signal through the ADC into the
system and (supposedly) to the DAC at channel 1. I was not able to see any difference on the scope when looking at channel one of the DAC.
I could not do a similar test to the ADC for comparison because it was not a SCSI connector inside the box.
I was able to see my signal when I connected the breakout board to the ADC input outside the PCIX box.
I then turned everything off, undid my troubleshooting rig, and closed the PCIX box up. I restarted again, and appear
to have all former functionality, so it seems I broke nothing new in my exploration.
I now suspect that it has something to do with the following errors in the code:
when I do a startoms after ssh'ing in, I get (among other things):
epicsThreadOnce0sd epicsMutexLock failed
Also, in the log file located at /cvs/cds/caltech/target/c2oms/log.txt the last three lines read:
Triggered the ADC
Net fail to fb0
DIAG RESET
Both of these things sound ominous, so I shall point them out to either Rob or the Russian, whoever I can get my hands on first. |
77
|
Mon Aug 11 15:32:00 2008 |
Stefan Ballmer | Electronics | General | no DAC output voltage when running OMS didgital system |
I tried to use the OMS digital system for a measurement, but I couldn't get the DAC output to work.
Symptoms:
- All EPICS buttons work, and digital read-backs make sense
- ADC works just fine
- DAC output voltage is zero, no matter what the filter module output says
- Watchdog is enabled (green) |
76
|
Mon Aug 11 14:15:57 2008 |
Dmass | Computing | CDS | ATF system |
Quote: | I made a preliminary generic model in simulink. This is stored on ws1 as /users/dmass/atf.mdl
I copied this to /cvs/cds/advLigo/src/epics/simLink/atf.mdl on machine oms.
I changed directories to /cvs/cds/advLigo and ran the make commands. Everything seemed to run, but there was no data coming through (fb0 seemed to not start up). I then found this post in the SUS_Lab elog by Sam. I ran the commands it said to run because of the 'broken nature of the DAQ code,' and everything seemed to work.
In the end, I ran the following sequence of commands on oms in the /cvs/cds/advLigo directory, and was able to get the frontend up and receiving data. The daq, however, is still not writing anything.
*make clean-atf atf install-atf
*rm /cvs/cds/caltech/chans/daq/C2ATF.ini (this step is needed because the DAQ install code isn't quite right at the time of this writing
*make install-daq-atf
*make install-screens-atf |
I talked to Alex, and he told me something about "your frames file is full, let me empty it and call you back," and he did so.
I did a reinstall of the daq
make unistall-daq-oms
and
make install-daq-oms
then I edited the .ini files with the channel names I wanted information about in
/cvs/cds/caltech/chans/daq/
C2OMS.ini is the file with the information I want in it.
oms.ini is the file which daqconfig interfaces with.
I edited the files via the daqconfig, and then overwrote the C2OMS.ini file with it, then data seemed to be recording and I could access it via dataviewer.
One concern/piece of information:
Apparently at some point they decided to change the convention on file suffixes for channel names. C2OMS.ini has a *_32768 tacked onto the end of
the channel name upon generation by the make scripts, whereas the oms.ini file has a *_DAQ in its place. This seemed to be the cause for some
of the noncooperation between the various sytems. Daqconfig likes the *_DAQ suffix, and this seems to jive with the data now writing, so we shall
go with that. |
75
|
Tue Aug 5 15:23:38 2008 |
Dmass | Computing | CDS | ATF system |
Quote: | I made a preliminary generic model in simulink. This is stored on ws1 as /users/dmass/atf.mdl
I copied this to /cvs/cds/advLigo/src/epics/simLink/atf.mdl on machine oms.
I changed directories to /cvs/cds/advLigo and ran the make commands. Everything seemed to run, but there was no data coming through (fb0 seemed to not start up). I then found this post in the SUS_Lab elog by Sam. I ran the commands it said to run because of the 'broken nature of the DAQ code,' and everything seemed to work.
In the end, I ran the following sequence of commands on oms in the /cvs/cds/advLigo directory, and was able to get the frontend up and receiving data. The daq, however, is still not writing anything.
*make clean-atf atf install-atf
*rm /cvs/cds/caltech/chans/daq/C2ATF.ini (this step is needed because the DAQ install code isn't quite right at the time of this writing
*make install-daq-atf
*make install-screens-atf |
The rm /cvs... command should be replaced by make uninstall-daq-atf. It does the removal and slightly more. |
74
|
Fri Aug 1 15:06:11 2008 |
Dmass | Computing | CDS | ATF system |
I made a preliminary generic model in simulink. This is stored on ws1 as /users/dmass/atf.mdl
I copied this to /cvs/cds/advLigo/src/epics/simLink/atf.mdl on machine oms.
I changed directories to /cvs/cds/advLigo and ran the make commands. Everything seemed to run, but there was no data coming through (fb0 seemed to not start up). I then found this post in the SUS_Lab elog by Sam. I ran the commands it said to run because of the 'broken nature of the DAQ code,' and everything seemed to work.
In the end, I ran the following sequence of commands on oms in the /cvs/cds/advLigo directory, and was able to get the frontend up and receiving data. The daq, however, is still not writing anything.
*make clean-atf atf install-atf
*rm /cvs/cds/caltech/chans/daq/C2ATF.ini (this step is needed because the DAQ install code isn't quite right at the time of this writing
*make install-daq-atf
*make install-screens-atf |
73
|
Fri Aug 1 14:52:27 2008 |
Dmass | Computing | DAQ | Daq not recording |
Quote: | There is still something wonky with the DAQ. Nothing appears to be writing at the moment, as I can only get 8 seconds of data from the playback in dataviewer. I may table this until I am trying to get the new system up and running. |
I abducted Rob from the 40m to look at this, and he thinks it has something to do with the fact that awgtpman (the test point manager) will not run. He beat at it with his kung-fu, but the computer would not yield. Ws1 does not yet know fear, but it shall soon learn. |
72
|
Tue Jul 29 00:58:36 2008 |
Masha | Laser | Fiber | first fiber noise measurement |
I'm attaching the fiber noise measurement from one mach zehnder channel
along with some noise sources. The total rms phase noise comes out to about
0.075 radians, which translates to a frequency noise of 20Hz rms.
I will take spectra at lower frequencies once the DAQ is back up so I don't
have to wait an hour for the spectrum analyzer. |
Attachment 1: MZnoise_sources0727.png
|
|
71
|
Tue Jul 29 00:47:37 2008 |
Masha | Laser | Fiber | mach zehnder setup |
As per dmass's prodding, I am putting up some pictures of what I'm doing.
Currently there is a working Mach Zehnder, but it needs to be improved.
Right now the fluctuations due to phase are about 15% of the total DC voltage,
where they were over 50% back at the 40m.On the list are more stable mounts
and better mode matching for fiber input and output with the laser. Also, we
will be putting in an AOM in the fiber arm. |
Attachment 1: setup1.JPG
|
|
Attachment 2: setup3.JPG
|
|
Attachment 3: setup4.JPG
|
|
70
|
Mon Jul 28 15:58:49 2008 |
Dmass | Computing | DAQ | Daq not recording |
There is still something wonky with the DAQ. Nothing appears to be writing at the moment, as I can only get 8 seconds of data from the playback in dataviewer. I may table this until I am trying to get the new system up and running. |
69
|
Fri Jul 25 17:07:34 2008 |
Dmass | Laser | PMC | Cavity Pole Measurement on the PMC |
Quote: | Measurement of the PMC cavity pole:
The following was done at powers of .234 W and .42 W.
I turned the laser power on the table down low (using a polarizer and a PBS to dump the majority of the beam).
I then used a function generator to input a triangle wave into the Piezo driver and sweep the cavity. This was done at a frequency of .04 Hz (for reasons dictated by dataviewer).
I found the HWHM (half width half max) of the transmitted intensity (when the DC bias was subtracted out) to be 0.00054 and 0.00053 free spectral ranges for .234 and .42 W, respectively.
I also noticed some nonlinearity in the PMC/Piezo as I scanned through more FSR's, and thus used the data from the bias voltages closest to zero.
I calculated the FSR frequency to be 713.8 MHz (in agreement to what we have on file for the PMC at the 40m).
Using the equality \delta L / L = \delta f / f, I get pole frequencies of 385 kHz and 378 kHz for .234 and .42 W, which is about a 2% agreement between measurements.
Again - the cavity pole was measured to be about 380 kHz to within a couple percent. - this differs significantly from the pole quoted for the 40m PMC at
http://www.ligo.caltech.edu/~ajw/PSLPMC.png (which is 488 kHz)
|
I looked back at the old data, and measured the V/cts ratio of the ADC (which at low frequency was 0.000622 V/ct). Using this, and the data from the above measurement of the cavity pole, the response of the PMC PZT is about 3.235 V/FSR.
Nota Bene:
The old data from the sweep of the PMC was taken by driving with a function generator, and tee-ing off the output of the generator to the ADC. However, I got the data from the output of one of the ADC channels, so it was post filtration. I am ALMOST POSITIVE that all gains were unity, but there is some small but finite possibility that one of them was not, and this would bebunk my number for Volts per free spectral range. |
68
|
Fri Jul 25 15:23:14 2008 |
Dmass | Laser | PSL | New Vertical PSL |
With the installation of the new "clear top" on the PSL (see picture), an interesting design feature has come to light.
There are a few noteworthy things about this feature:
-We have a vertically oriented window after the NPRO.
-The vertically oriented window shoots some beam up.
-The new clear top appears to be transparent to IR.
-The HEPA filter is right above the window.
Possible worries:
-Light leaking back into the NPRO.
-Dumping beam onto the HEPA filter might not be the best thing.
It might be a good idea to do a proper dumping of that beam. I am unsure as to whether it would be better to do so inside or outside the enclosure. |
Attachment 1: NPRO.png
|
|
Attachment 2: HEPA.png
|
|
67
|
Fri Jul 25 14:59:24 2008 |
Dmass | Computing | fubar | digital control broken. |
Quote: | I tried to restore the hacked up oms control system via a series of escalating measures.
a) I ssh'ed into the FE computer (oms) and did "startoms"
-that seemed to run, but there were errors in the status window of screen "C2OMS_GDS_TP.adl" next to the red light FB0, of type 0x2001 if I recall correctly
b) I then tried to remake the oms FE code via the commands in Tobins elog entry
- make oms
- make install-daq-oms
- make install-oms
- make install-screens-oms
Nothing upon restarting the FE. No more error messages from the FB0 button now. possibly more broken than before.
c) Stefan and I then did a "sudo reboot" on oms, still nothing.
We found as the last line in the log file
/cvs/cds/caltech/target/c2oms/log.txt
"DAQ init failed -- exiting"
It appears that the daq no longer starts. I have sent Alex an email, and have tried calling him several times (have not gotten a pickup).
Daqconfig is also brokenish on the non FE computer (ws1). It gives "Error: can't read "y": no such variable" on startup, and will not display anything.
I will be working on getting Rich Abbott's PDH box up and running to lock the PMC while I twiddle my thumbs on the digital stuff. |
With the assistance of a benevolent Rob, the above problem was identified and fixed.
While looking at the logfile for C2oms, we found the script omsfe.rtl was crapping out because there were no daq channels set to acquire. We fixed this by manually setting one of the channels to acquire in C2oms.ini and uncommenting it. We then dropkicked the framebuilder by telneting in via
> telnet fb0 8087
and doing a shutdown. This should be equivalent to clicking the DAQ Reload button on the oms FE Diagnostic medm screens. This fixed the problem of the daq not starting, and we could get realtime data via dataviewer. To start recording it in the daq again, we just need to edit "C2oms.ini" to reflect which channels we are interested in recording.
It appears we can still not toggle channels to acquire via the daqconfig gui. I will worry about that more when I have a version control screens here that I am happy with, and harrass Alex.
We also found something interesting when we tried to interact with the system via DTT. It seemed that the system wanted there to be a link tpchanm1 pointing to tpchanC2, but the creation and subsequent annhilation of this link seemed to not effect the system. When we looked closer, the file we expected to define all the test points, /cvs/cds/caltech/chans/param/tpchn_C2.par, was full of objects for the atf system. The file that the system appears to actually get its definitions of these from was /cvs/cds/advLigo/build/omsepics/oms.par.
Rob says that this was an indication that something in the series of "make ___" commands is broken, as they should have overwritten this file at some point. |
66
|
Wed Jul 23 12:04:01 2008 |
Dmass | Electronics | General | Universal PDH box |
Quote: | I just picked up the "Universal PDH Box" prototype from Rich Abbott.
A few wrinkles: One of the voltage regulators was resonating at 12.3 kHz, producing an annoying high pitched whine, so Rich added a beefy capacitor in parallel with the original SM component.
There was possible shorting of the bottom of the circuit board to the case. I added some rubber feet to the back of the board, and covered the inside of the case in some kapton tape. I will test again later (possibly when we get the NIM crate down here).
We (I) need to get a NIM crate cooking to populate with useful things like this.
Rich also advised triple checking that the board is powered correctly from the NIM supply. Will do this when I get a crate.
We did some rudimentary testing of the box in Wilson house, and it seems to be functioning as expected.
Also should make sure that we are only AC coupled to the photodiode when we finally do hook one up to the front.
Thorough (aka large) picture attached.
Also all pertinent documentation attached below picture in a pdf. |
I measured the transfer function of the PDH Box between the "Servo Input" and "Piezo Drive Output" terminals on the front of the box.
I used the 4395, and for 10 Hz - 200 kHz I used a BNC T-splitter for splitting the drive signal. for 1kHz - 35MHz I used a minicircuits ZMSC-2-2 splitter (which is rated in that frequency band).
Attached are both transfer functions for posterity. I shall try to lock the PMC with this next. |
Attachment 1: PDH_LF.png
|
|
Attachment 2: PDH_HF.png
|
|
65
|
Tue Jul 22 18:43:06 2008 |
Dmass | Computing | fubar | digital control broken. |
I tried to restore the hacked up oms control system via a series of escalating measures.
a) I ssh'ed into the FE computer (oms) and did "startoms"
-that seemed to run, but there were errors in the status window of screen "C2OMS_GDS_TP.adl" next to the red light FB0, of type 0x2001 if I recall correctly
b) I then tried to remake the oms FE code via the commands in Tobins elog entry
- make oms
- make install-daq-oms
- make install-oms
- make install-screens-oms
Nothing upon restarting the FE. No more error messages from the FB0 button now. possibly more broken than before.
c) Stefan and I then did a "sudo reboot" on oms, still nothing.
We found as the last line in the log file
/cvs/cds/caltech/target/c2oms/log.txt
"DAQ init failed -- exiting"
It appears that the daq no longer starts. I have sent Alex an email, and have tried calling him several times (have not gotten a pickup).
Daqconfig is also brokenish on the non FE computer (ws1). It gives "Error: can't read "y": no such variable" on startup, and will not display anything.
I will be working on getting Rich Abbott's PDH box up and running to lock the PMC while I twiddle my thumbs on the digital stuff. |
64
|
Fri Jul 18 16:14:28 2008 |
Dmass | Electronics | General | Universal PDH box |
I just picked up the "Universal PDH Box" prototype from Rich Abbott.
A few wrinkles: One of the voltage regulators was resonating at 12.3 kHz, producing an annoying high pitched whine, so Rich added a beefy capacitor in parallel with the original SM component.
There was possible shorting of the bottom of the circuit board to the case. I added some rubber feet to the back of the board, and covered the inside of the case in some kapton tape. I will test again later (possibly when we get the NIM crate down here).
We (I) need to get a NIM crate cooking to populate with useful things like this.
Rich also advised triple checking that the board is powered correctly from the NIM supply. Will do this when I get a crate.
We did some rudimentary testing of the box in Wilson house, and it seems to be functioning as expected.
Also should make sure that we are only AC coupled to the photodiode when we finally do hook one up to the front.
Thorough (aka large) picture attached.
Also all pertinent documentation attached below picture in a pdf. |
Attachment 1: PDH_User_Manual1.pdf
|
|
Attachment 2: UniPDH.png
|
|
63
|
Thu Jul 17 16:44:27 2008 |
Dmass | Laser | PSL | Laser is off |
Leaving laser off while I hack at screens downstairs. Feel free to turn on for anything.
I also turned the ISS drive box off. |
62
|
Wed Jul 16 11:47:18 2008 |
Dmass | Computing | General | Front End Machine Namechange |
The front end machine in the Bridge Basement has been renamed "atf" from "oms".
I changed /etc/sysconfig/network on oms(now atf) and
/etc/hosts on ws1 (the dual head control station)
by replacing the "oms" tag with "atf"
When you ssh into atf from ws1, it still says you are logged in as "controls@oms". There appears to be more to change.
Disambiguation of the digital locking system to follow. |
61
|
Mon Jul 14 16:01:45 2008 |
Stefan Ballmer | Electronics | ISS | H1 ISS spare tested |
I tested the modified H1 ISS spare using a post-PMC sensing.
I got the expected performance in-loop (about 2e-8/Hz at 100Hz),
but the out-of-loop diode had additional acoustic noise. But that's not due to the electronics.
Also, for this test, I again locked the PMC using the digital system. |
60
|
Fri Jul 11 16:20:21 2008 |
Stefan Ballmer | Electronics | | PMC |
Since the digital system was in an ambigous state I locked the PMC using a SR560.
I did this so I can use the lab setup to test the newly modified Hanford ISS spare. |
59
|
Sun Jun 1 18:52:49 2008 |
tobin | Computing | General | borkspace compiling |
Following Alex's suggestions, I fixed the front-end problems we were encountering the other day.
To get the shared libraries stuff to work, I inspected the contents of /etc/ld.so.conf.d and
found it all looking fine. I ran /sbin/ldconfig, which updates some shared library index
somewhere, which seemed to fix the problem with the EPICS server finding its libraries. To create
the /rtl_mem_atf file, I edited /etc/rc.local, adding "atf" to the list of subsystems. I believe
that everything now works to compile and run the ATF front-end system.
As a reminder, some useful commands:
cvs update refreshes the software distribution from CDS
make atf recompiles the front-end system
make install-atf installs new front-end binary and scripts
make install-daq-atf installs DAQ channel stuff to framebuilder
make install-screens-atf makes new generic MEDM screens
killatf Stop the front end code
startatf Stop and then start the front-end-code
If you want to change the name of the computer itself, I think you just
need to (1) edit /etc/sysconfig/network and change the hostname; and (2)
etc /etc/hosts on all the machines from which you'll be connecting to the
front-end machine (at the moment, just ws1?).
[fixed command syntax - 7/22/08 DYM] |
58
|
Thu May 29 20:00:48 2008 |
tobin | Computing | Fugly | borkspace compiling |
Dmass and I tried to compile and install a new system on the borkspace machine here.
We ran into a few difficulties.
First, the startup script for the EPICS server can't find the channel access library.
A hack to fix this is to enter this command in the terminal before running the start
script:
export LD_LIBRARY_PATH=/opt~/epics-3.14.7-x86_64/base-3.14.7/lib/linux-x86
With this, the OMS system front-end runs fine. But we can't start our new system,
ATF. The ATF front-end dies with the error:
[controls@oms advLigo]$ cat /cvs/cds/caltech/target/c2atf/log.txt
cpu clock 2412402
open failed for write on /rtl_epics (45)
I looked in the controller.c code and it appears that the front-end code wants there
to be a special file /rtl_mem_atf. I made one of these using both a hard link, and
then via mknod ("sudo mknod rtl_mem_atf u 150 132"), mirroring the existing rtl_mem_oms,
but neither attempt worked. Will ask Alex for help. |
57
|
Tue May 20 15:55:31 2008 |
Dmass | Laser | EOM | EOM Phase lag |
The attached transfer function was measured with the experimental setup outlined in entry 53. The first transfer function (Raw Data) is the output of the 4395 when I calibrated away the cabling as described in the drawing in entry 53. The high frequency junk is pickup from something; it shows up with the photodiode blocked.
The second transfer function is the first one with the following model divided out:
model=@(f) (i*f).*(1./(1+i*f/380e3)).*exp(-i*2*pi*f*60e-9)
The zero at zero is due to the EOM modulating in phase, not frequency.
The time delay was generated by roughly fitting to the data (to get a white phase response)
The pole at 380 kHz is the cavity pole of the PMC (experimentally determined in entry 56)
The measured time delay based on propagation of cables/free space is 41 ns, so there is an additional time delay of 19 ns +/- 10 ns in our model.
This corresponds to a phase shift of 0.7 degrees at 100 kHz
As we can see from the transfer function - there is clearly more going on than our model takes into account (probably a number of poles). Adding poles would put a tighter constraint on the time delay, which is fine/good.
Quote: | I used the experimental setup below to measure the transfer function between the EOM (which is before the amplifier) and the reflections off the PMC. The light blue line is what I shorted to each other when I used the 4395's calibrate feature to correct for the cabling/rf components.
The "modulation summation box" is shown below the schematic. I will update this later when I figure out what the components are. |
|
Attachment 1: RawTF.png
|
|
Attachment 2: corrtf.png
|
|
56
|
Mon May 19 17:28:44 2008 |
Dmass | Laser | PMC | Cavity Pole Measurement on the PMC |
Measurement of the PMC cavity pole:
The following was done at powers of .234 W and .42 W.
I turned the laser power on the table down low (using a polarizer and a PBS to dump the majority of the beam).
I then used a function generator to input a triangle wave into the Piezo driver and sweep the cavity. This was done at a frequency of .04 Hz (for reasons dictated by dataviewer).
I found the HWHM (half width half max) of the transmitted intensity (when the DC bias was subtracted out) to be 0.00054 and 0.00053 free spectral ranges for .234 and .42 W, respectively.
I also noticed some nonlinearity in the PMC/Piezo as I scanned through more FSR's, and thus used the data from the bias voltages closest to zero.
I calculated the FSR frequency to be 713.8 MHz (in agreement to what we have on file for the PMC at the 40m).
Using the equality \delta L / L = \delta f / f, I get pole frequencies of 385 kHz and 378 kHz for .234 and .42 W, which is about a 2% agreement between measurements.
Again - the cavity pole was measured to be about 380 kHz to within a couple percent. - this differs significantly from the pole quoted for the 40m PMC at
http://www.ligo.caltech.edu/~ajw/PSLPMC.png (which is 488 kHz)
|
55
|
Tue May 13 12:37:31 2008 |
Dmass | Laser | Fiber | New fiber |
We got a box of fiber from Boston (Marie Woods). It is now on the optics table in the bridge basement.
The fiber has some writing on it. It says printed on the fiber in ink:
"OZ OPTICS LTD BC# 2887" and gives their tel # as 613-831-0981.
There is also a label on the fiber that says:
"PMJ-3AF,3AF-1064-6/125-3-60-1
SN: T872391-01"
Picture below: |
Attachment 1: 00011.png
|
|
54
|
Wed May 7 04:44:31 2008 |
Dmass | Computing | General | DTT and Fast Channels fixed |
The playback in dataviewer on the fast channels was semi broken (would not display full data on a playback, but could do all else).
Also, DTT was giving an error "failed to start diagnostics kernel" whenever we attempted to launch it.
I talked to Alex and he managed to fix both of these problems. |
53
|
Tue May 6 18:18:59 2008 |
Dmass | Laser | EOM | EOM Phase lag |
I used the experimental setup below to measure the transfer function between the EOM (which is before the amplifier) and the reflections off the PMC. The light blue line is what I shorted to each other when I used the 4395's calibrate feature to correct for the cabling/rf components.
The "modulation summation box" is shown below the schematic. I will update this later when I figure out what the components are. |
Attachment 1: TF_setup1.png
|
|
Attachment 2: MODSUM.png
|
|
52
|
Mon May 5 20:32:50 2008 |
Dmass | Computing | General | MEDM, FB, FE, Etc |
I restarted the oms box in the rack, and was in the process of trying to get new medm control screens up with Rob. I now have neither the old mislabelled oms screens, nor the new soon to be beautiful screens, as well as no way to finish my measurement of the PMC cavity pole until I get those back up and running.
Will be reading
http://www.ligo.caltech.edu/%7Eaivanov/daq_handbook.html
and if that fails,
Will bug Alex or Stefan tomorrow. |
51
|
Fri May 2 22:54:44 2008 |
Dmass | Computing | | Screens |
I made a simulink model and put it here: /cvs/cds/advLigo/src/epics/simLink/atf.mdl
as instructed by the rolf wiki:
http://lhocds.ligo-wa.caltech.edu:8000/40m/WikiOfRolfCDS?highlight=%28rolf%29
I then changed to the advLigo directory on oms, and "make atf" seemed to run successfully. Before finishing the process, I got the following error when dealing with c2atfepics:
[controls@oms c2atfepics]$ cat iocC2.log
dbLoadDatabase "dbd/a.dbd"
dbLoadDatabase "dbd/atf.dbd"
dbLoadDatabase "dbd/get_local_time.dbd"
registerRecordDeviceDriver(pdbbase)
dbLoadRecords "db/C2/local_time.db"
dbLoadRecords "db/C2/atf1.db"
iocInit
Starting iocInit
############################################################################
### EPICS IOC CORE built on Aug 10 2005
### EPICS R3.14.7 $R3-14-7$ $2004/12/06 22:31:52$
############################################################################
iocInit: All initialization complete
seq &atf,("ifo=C2, site=caltech, sys=ATF, sysnum= 10")
SEQ Version 2.0.10: Wed Aug 10 23:41:07 2005
cas warning: Configured TCP port was unavailable.
cas warning: Using dynamically assigned TCP port 46391,
cas warning: but now two or more servers share the same UDP port.
cas warning: Depending on your IP kernel this server may not be
cas warning: reachable with UDP unicast (a host's IP in EPICS_CA_ADDR_LIST)
Couldn't open `/rtl_mem_atf' read/write
open("rtl_epics"): No such file or directory |
50
|
Fri May 2 22:00:17 2008 |
Dmass | Computing | | Daq Channels |
I activated the channels C2:OMS-SUS_T3_ACT_OUT and C2:OMS-SUS_LEFT_IN1 in daqconfig, set both to acquire, and saved them to cvs/cds/caltech/chans/daq/C2OMS.ini
Real screens and naming conventions to follow.
[Edit: I do not know how to disable the auto smileys, so : O is seen as ] |
49
|
Fri May 2 21:39:13 2008 |
Dmass | Laser | PSL | Laser power on table lowered |
I have changed the polarizer before the polarizing beam splitter to minimize output power.
I did this to measure the PMC pole at as low a power as possible.
This will make power trends (which are picked off after the PBS) look VERY low for a bit. |
48
|
Fri May 2 21:09:26 2008 |
Dmass | Laser | General | Strange Noise |
We may want to set up a microphone and see if the noise is correlated to the sharp power fluctuations in the laser.
Quote: | Something on the table or under it is periodically making a squeeking noise. Stefan claims this is not new though neither of us can localize it (it's < 1sec duration). Unsure as to what is the offending noisemaker.
[Update - the noise happens when the laser is off, comes from side of table (possibly beneath) opposite laser] |
|
47
|
Fri Apr 25 13:33:35 2008 |
Dmass | Laser | PSL | power drift |
Here is a trend of the power drift taken over 10 days. The units are in counts produced by a photodiode stationed at a lower power pickoff of the beam.
The first plot is a full 10 days - you can see the times during which the laser was off where the counts go to zero.
The second plot is the full data over 10 hours, which seems to imply that the sharp increases in power are real.
Negative values correspond to an increase in power. |
Attachment 1: trend2.png
|
|
Attachment 2: trend3.png
|
|
46
|
Tue Apr 22 20:07:50 2008 |
Dmass | Laser | PSL | Profile of 35W PSL |
The total power seemed to be down about 10%, and that lower number was divided between two beams at the output. They opened up the amplifier and realigned things within it, which seems to have fixed the multibeam/lower power we were encountering before.
Quote: |
Quote: | I scanned the beam on the 35W laser again (for the first time since it was fixed by our German friends during the LSC meeting).
|
What did they fix? And how? |
|
45
|
Tue Apr 22 16:25:24 2008 |
rob | Laser | PSL | Profile of 35W PSL |
Quote: | I scanned the beam on the 35W laser again (for the first time since it was fixed by our German friends during the LSC meeting).
|
What did they fix? And how? |
44
|
Fri Apr 11 17:15:49 2008 |
Dmass | Laser | EOM | Transfer Function |
We traded network analyzers and now can go below 100kHz in our measurement. The following transfer function was taken with an HP 4395A.
The previous measurements were done with the PMC not locked to the laser, so the measurement was repeated using RF Reflection locking (sidebands @ 35MHz).
Another important thing we discovered was that the frame of the table is ungrounded, which led to significant noise leaking into the error signal when we let a BNC cable touch it. Schematic of experimental set up to follow.
Not all of the cables/electronics were calibrated out in the measurement below. |
Attachment 1: EOM_PMC_tf.png
|
|
43
|
Fri Apr 11 16:37:03 2008 |
Stefan Ballmer | Computing | General | Set up CDS wireless router / controlling from nokia handheld works |
I set up a CDS wireless router:
IP: 131.215.113.2
SSID: cdsrana
Unname & PWD written on router
It has WEP and MAC-filtering enabled.
With that I was able to use the little Nokia handheld to control the digital system.
Example:
- start x-terminal
- 'cd'
- 'ws1'
(this just starts ssh to ws1)
- startSlider
(this bring up the offset slider for the PZT control loop), or
- startmedm
(this starts the OMS main screen)
We have to make medm screens that work for the little Nokia screen.... |
42
|
Mon Apr 7 16:43:15 2008 |
Dmass | Laser | PSL | Profile of 35W PSL |
I scanned the beam on the 35W laser again (for the first time since it was fixed by our German friends during the LSC meeting).
A picture of the setup is below.
Images of the profile, 2d and 3d are below as .bmp, and the generating data for these is in Apr7_08scan.asc (zipped text)
This profile is significantly cleaner than the one measured in entry 31 of the elog. |
Attachment 1: scancap.png
|
|
Attachment 2: 3dbeamprof.bmp
|
Attachment 3: flatprof.bmp
|
Attachment 4: Apr7_08scan.asc.zip
|
41
|
Thu Apr 3 17:56:35 2008 |
Dmass | Laser | EOM | Transfer Function |
Transfer function part 2:
I repeated the measurement with impedence matching where I could (not inside the laser box), and did some CRUDE fitting with matlab.
I also herby acknowledge that png > jpg.
The tau is the time delay which makes the green curve on the phase plot coincide with the data at high frequency.
No locking was used. The 00 mode was reached for short periods of time through input to the laser crystal temperature during which time the measurement was taken.
-Calibration-
The input to the EOM was removed and connected to the PD output and calibrated out with the network analyzer.
The analyzer used to do this measurement has now been traded for a HP 4395A (which goes down to 10Hz). |
Attachment 1: EOMPMCTF.png
|
|
40
|
Tue Apr 1 20:50:37 2008 |
Dmass | Laser | EOM | Transfer Function |
(With Stefan's help) I took the transfer function between the EOM of the PSL, and the Photodiode on the PMC's reflected light. We looked at the lower frequency as well with another analyzer (up to 50 MHz).
The very low frequency features seem to be junk, and we can see what looks like ~f in the low frequency transfer function.
In the higher frequency range, there seem to be resonances corresponding to a FSR of about 22 MHz (long.) starting around 11 MHz. The low end seems to behave as expected, and go as frequency. (Note that the units on the non-log bottom plot are in x10^7 Hz) |
Attachment 1: EOM_PD_xsfer.jpg
|
|
Attachment 2: HFlogxsfer.jpg
|
|
Attachment 3: HFxsfer.jpg
|
|
39
|
Thu Mar 27 13:25:38 2008 |
Stefan Ballmer | Laser | ISS | Modifications to ISS (D080149-00-C) |
|
Attachment 1: D080149-00-C.pdf
|
|
38
|
Fri Mar 7 15:33:14 2008 |
Dmass | Laser | EOM | EOM shipped |
I packaged up and shipped the EOM back to UF. The translation stage, also theirs, was sent with it. Everything is marked c/o Guido Mueller. |
37
|
Tue Mar 4 16:00:28 2008 |
Dmass | Laser | PSL | ISS |
I moved the (non logging) ISS PD after the PMC, and changed the mirror/camera setup some.
Setup post PMC should now have ~5% of the power going to the camera, with the rest going to a BS for the ISS photodiodes. |
36
|
Mon Mar 3 01:21:19 2008 |
rana | Laser | MOPA | power drift |
This plot shows the power drift over the last few days.
The SIDE_OUT16 channel is the one. The y scale is zoomed in here but
zero corresponds to zero power. The faster spikes have a ~45 minute time scale.
No idea what this is yet. |
Attachment 1: test.png
|
|
Attachment 2: test2.png
|
|
35
|
Wed Feb 27 17:56:45 2008 |
Stefan Ballmer | Electronics | General | |
The 2nd HP 6209B 320V power supply initially didn't work because one of its backplane connectors had fallen off.
It has been fixed, and is now on the shelf, ready to be used. |